# A Catalog of $\sim 146,000$ Radio Galaxies up to $z\sim 5$

### v1

@@ 5.2 High-$z$ Candidates with $z\geq 3$
As introduced in § 5.1, the USS characterized by $\alpha<-1.3$ is an efficient technique to select HzRGs. However, in addition to these USS sources, many HzRGs have been discovered with relatively flat spectral index (e.g., ${\alpha}_{\mathrm{1.4GHz}}^{\mathrm{325MHz}}=-0.75$ for J163912.11+405236.5 at $z=4.88$, Jarvis et al. 2009; ${\alpha}_{\mathrm{1.4GHz}}^{\mathrm{325MHz}}=-1.1$ for HSC J083913.17+011308.1 at $z=4.72$, Yamashita et al. 2020; see also Miley & De Breuck 2008). As a consequence, a sole reliance on USS could lead to a bias in the statistical physical properties of the HzRGs since the diverse nature could be over-represented by USS.
An alternative approach to select HzRGs, which can circumvent the selection bias arising from the uncertain real spectral index distributions, is the Lyman break technique commonly adopted in the search for high-$z$ galaxies. The effectiveness of this technique in discovering HzRGs has been proven by Capetti et al. (2025), who identified 39 $g-$dropout HzRGs expected to be at $4.53<z<$ 4.5 over an area of $\sim 2,000\mathrm{\ deg^{2}}$, and by Yamamoto et al. (2025), who constructed a photometric sample of 146 HzRGs at $z\sim 4$ within a total survey area of $\sim 560\mathrm{\ deg^{2}}$.
The Lyman break color selection technique selects sources that exhibit a clear Lyman break and blue UV continuum in their optical SEDs covered by broadband (ugriz) photometry. Following Hildebrandt et al. (2009) that have used a similar filter set, g-dropouts ($53<z<$ 5) are selected according to
$g-r>1.0,$ (6)
$r-i<1.0,$
$g-r>1.5(r-i)+0.8,$
and r-dropouts ($64<z<$ 6) are extracted adopting
$r-i>1.2,$ (7)
$i-z<0.7,$
$r-i>1.5(i-z)+1.0.$
@@
We show g- and r-dropout candidates in the upper and lower panels of Figure 17, respectively. The above criteria produce 4,049 valid sources for the r-dropouts, and only 12 of them can be at $z\sim 5$, which is reasonable considering the limited depth of UNIONS. Of ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS sources with reliable BPZ photo-$z$, there are only 2 at $z>4.5$, comparable with the r-dropout selections. The number of g-dropout-selected sources is $477$, which exceeds the number counts of 53 expected from reliable $4.53.5<{z}_{\mathrm{photo}}<$ 4.5 estimates by almost one order of magnitude.
This potential excess in the number counts of g-dropouts – if they are not bona fide sources at $z\sim 4$ – could be explained by the contamination of foreground sources such as red elliptical galaxies at intermediate redshifts, whose ${D}_{\mathrm{4000}}$ break due to the absorption of old stellar populations mimic the Lyman break feature. The photometric errors of these foreground contaminations satisfy the selection criteria of the g-dropout (e.g., Ono et al., 2018). For Subaru HSC, which has comparable optical number counts and similar efficient broadband wavelengths to those of UNIONS, Ono et al. (2018) found the contamination rate in the $i-$band magnitude bin of 22.5 can reach $\sim 68\%$ and decreases to $\sim 41\%$ at ${i}_{\mathrm{ABmag}}=23.1$ and to $\sim 35\%$ at ${i}_{\mathrm{ABmag}}=23.7$ for g-dropouts. Therefore, even if we only select UNIONS galaxies at the faint end of $i-$band (${i}_{\mathrm{AB}}>23$), the contamination problem can still persist. Similarly, for $r-$dropouts, Ono et al. (2018) reported contaminate rates of approximately $23\%$ at ${z}_{\mathrm{ABmag}}=22.3$ and around $59\%$ at ${z}_{\mathrm{ABmag}}=23.5$, implying that merely $<5$ $r-$dropouts could be true positives if applying these estimates to UNIONS-VLASS. Nonetheless, under the optimistic assumption of HSC-level contamination rates, our UNIONS-VLASS radio galaxy catalog is expected to yield more than 200 HzRGs at $z\sim 4$, offering strong potential to significantly enhance the statistical robustness of studies on HzRGs.
@@ 5.3 Caveat to the Catalog
The effect of applying $\mathtt{NN\_dist>30^{\prime\prime}}$ to VLASS2 is clearly reflected in the projected linear size, that is, the majority of our RGs are compact (${l}_{\mathrm{s}}\leq 20$ kpc), which means our UNIONS-VLASS RGs have the tendency to be young and evolving populations. This also leads to a potential bias of our HzRG populations. As they are often selected by the USS feature in previous studies, there are only a few of them showing USS at $z>2$ amongst $15,906$ UNIONS-VLASS-FIRST sources with valid ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$. This global trend of being compact for our RGs is more a result of the limited search radius of $1.8^{\prime\prime}$, As shown in Figure 2, in $\mathtt{NN\_dist<30^{\prime\prime}}$, there is no concentration of radio sources with small offsets ($\leq 1^{\prime\prime}$) from their optical counterparts. To maintain the current cleanliness of UNIONS-VLASS catalog, even if $\mathtt{NN\_dist<30^{\prime\prime}}$ is not adopted, we do not expect a significant increase in the number counts of bright and extended radio sources. We, thereby, remind the readers to be cautious about the intrinsic and unresolved bias arising from the design of this work. Nonetheless, the vast size of this catalog still ensures statistical studies of radio AGNs across various evolution stages, offering valuable insights into the connections between radio AGNs and their host galaxies.
@@ 6 Summary
We presented radio color-color (${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}-{\alpha}_{\mathrm{3000MHz}}^{\mathrm{1400MHz}}$) based on LoTSS, FIRST, and VLASS observations, and further plotted spectral indices as a function of redshift and the projected linear size of the synchrotron-emitting volume. We found a great diversity of radio classifications, including Compact-Steep-Spectrum and Peaked-Spectrum sources that are likely young radio galaxies, Ultra-Steep-Spectrum sources which may have significantly evolved, and inverted and upturned spectrum that may be linked to transient sources. There are $\sim 10,000$ additional radio galaxies corresponding to more than one optical counterpart. These sources are merger candidates that host radio galaxies. The radio classifications/RLAGN evolution and merger-RLAGN connections will be further studied in future works.
@@ Appendix A Radio Morphology of Sources with $\mathtt{NN\_dist}<30^{\prime\prime}$
We randomly choose 36 VLASS radio sources with $\mathtt{NN\_dist}<30^{\prime\prime}$ and show their Epoch 2 QL images in
@@ Appendix C UNIONS-VLASS r-band Source Counts for RL and RQAGNs
Figure 20: Optical $r-$band source counts for 63,019 UNIONS-VLASS radio galaxies (circles). In both panels, squares represent RL and crosses represent RQ populations. Left: the UNIONS-VLASS sources are divided into RL and RQAGNs based on the radio loudness, where RLAGNs have $\mathcal{R}_{\mathrm{obs}}>1$. Right: RL and RQAGNs are separated via the 1.4 GHz spectral luminosity, where RLAGNs have ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$.
@@
We further divide RL and RQ populations based on their 1.4 GHz spectral luminosities (see § 4.3 for details). We see that RQAGNs are still the dominant populations for galaxies at ${r}_{\mathrm{ABmag}}<18.5$. However, at ${r}_{\mathrm{ABmag}}>20$, the source counts of RQAGNs do not decrease as steeply as the trend based on $\mathcal{R}_{\mathrm{obs}}$, which is similar to that of radio quasars (see, e.g, Yamashita et al. 2018). With the increasing $r-$band magnitude, there is a lower possibility of finding RQAGNs compared to RL populations. In summary, from this aspect, both RL and RQAGNs are hosted by a large variety of galaxy populations.
@@ Appendix D Radio-loud AGNs
Star formation can contribute to the non-thermal emission, and rigorously $\log\mathcal{R}_{\mathrm{obs}}>2$ may be adopted to exclude star-forming/starburst galaxies that mimic RLAGNs (e.g., Chiaberge & Marconi, 2011). This reduces our RLAGNs to 109,901 ($\sim 77.0\%$). The spectral luminosity of free-free emission attributed to star formation can be quantified by (Condon, 1992):
$\left({\frac{{L}_{\mathrm{NT}}}{\mathrm{\ W\ Hz^{-1}}}}\right)\sim 5.5\times 10^{20}\left({\frac{\nu}{\mathrm{GHz}}}\right)^{-0.1}\left[\frac{\mathrm{SFR(\mathit{{M}_{\mathrm{star}}}\geq\mathrm{M_{\odot}})}}{\mathrm{M_{\odot}}\ \rm yr^{-1}}\right],$ (D1)
@@
$\left({\frac{{L}_{\mathrm{NT}}}{\mathrm{\ W\ Hz^{-1}}}}\right)\sim 5.3\times 10^{21}\left({\frac{\nu}{\mathrm{GHz}}}\right)^{-0\alpha}\left[\frac{\mathrm{SFR(\mathit{{M}_{\mathrm{star}}}\geq\mathrm{M_{\odot}})}}{\mathrm{M_{\odot}}\ \rm yr^{-1}}\right].$ (D2)
Assuming an upper limit of $\mathrm{SFR}=300\ \mathrm{M_{\odot}}\ \rm yr^{-1}$ for star-forming main sequence galaxies at $z\gtrsim 2$ (Popesso et al., 2023), the corresponding luminosities are ${L}_{\mathrm{T,SF}}\sim 1.5\times 10^{23}\mathrm{\ W\ Hz^{-1}}$ and ${L}_{\mathrm{NT,SF}}\sim 1.3\times 10^{24}\mathrm{\ W\ Hz^{-1}}$. In Figure 13 for ${L}_{\mathrm{1.4GHz}}$ and ${L}_{\mathrm{3GHz}}$ as a function of $z$, this ${L}_{\mathrm{NT,SF}}$ only selects radio galaxies at $z\lesssim 1$ where the host galaxies should have much smaller values of SFR compared to high-z sources even if they are rarely observed starburst systems. The real contributions of star-forming activities to the observed synchrotron radiation should be insignificant ($\lesssim 1.5\times 10^{22}\mathrm{\ W\ Hz^{-1}}$ assuming $\mathrm{SFR}=40\ \mathrm{M_{\odot}}\ \rm yr^{-1}$ at $z\sim 0.5$). Therefore, $\mathcal{R}_{\mathrm{obs}}>1$ can be safely adopted to select RLAGNs for the UNIONS-VLASS sample.
We see correlations between ${L}_{\mathrm{1.4GHz}}-\mathcal{R}_{\mathrm{obs}}$ and $\mathcal{R}_{\mathrm{obs}}-{z}_{\mathrm{photo}}$, and these correlations are likely a propagation of the correlation between ${i}_{\mathrm{ABmag}}-{z}_{\mathrm{photo}}$ shown in the left panel of Figure 21 that reflects the cosmological surface brightness dimming. As a result, high-$z$ sources have higher possibilities of being classified as RLAGNs because of their optical faintness. In the right panel of Figure 21, we plot absolute radio magnitude $M_{t}$ as a function of the observed radio loudness $\mathcal{R}_{\mathrm{obs}}$, where $M_{t}$ is defined as (Ivezić et al., 2002)
@@
Figure 21: Left: UNIONS $r-$band photometry as a function of the photometric redshift. Right: Absolute radio magnitude as a function of the observed radio loudness, where the dashed line indicates the RL/RQ demarcation of $M_{t}=-25.9$. The golden dots represent RLAGNs based on $\mathcal{R}_{\mathrm{obs}}>1$ and the violet ones are RQAGNs defined by $\mathcal{R}_{\mathrm{obs}}<1$.
@@ Appendix E UNIONS2000 versus UNIONS5000 for i-band Photometry
We compare the distributions of the $r-i$ color between UNIONS2000 and UNIONS5000 for 76,454 UNIONS-VLASS radio galaxies produced in both data releases in Figure 22. The median difference between ${r}_{\mathrm{ABmag}}$ and ${i}_{\mathrm{ABmag}}$ is 0.44 in both UNIONS2000 and UNIONS5000. However, in comparison with UNIONS2000, the distribution of $r-i$ in UNIONS5000 is more convergent, which is positively attributed to that the UNIONS5000 has been additionally processed using Pan-STARRS DR3 $i-$band data.
@@ Appendix F Catalog columns
We show delivered columns of our UNIONS-VLASS radio galaxy catalog in Table LABEL:tab:_catalog_columns, including VLASS whose radio source is set as the reference at 3 GHz, UNIONS (optical; ugriz), SDSS (optical; ugriz), Pan-STARRS (optical to NIR; grizy), AllWISE (NIR to MIR; 3.4, 4.6, 12, and 22 $µm\micron$m), unWISE (NIR; 3.4 and 4.6 $µm\micron$m), FIRST (1.4 GHz), and LoTSS (144 MHz), as well as properties calculated in this work.
@@
Columns for VLASS (Gordon et al., 2021)
@@
Columns for UNIONS (Gwyn et al., 2025)
@@
Columns for SDSS DR16 (Ahumada et al., 2020)
@@
1$=$bad; 2$=$barely acceptable; 3$=$good (field_quality) (13)
@@
Columns for Pan-STARRS DR1 (Chambers et al., 2016)
@@
Columns for AllWISE (Cutri et al., 2021)
@@
Columns for unWISE (Schlafly et al., 2019)
@@
Columns for FIRST (Helfand et al., 2015)
@@
Columns for LoTSS DR2 (Shimwell et al., 2022)
@@
flag_z_photo int -1$=$no valid photo-$z$;
[…]

### v2

@@ 5.2 High-$z$ Candidates with $z\geq 3$
As introduced in § 5.1, the USS characterized by $\alpha<-1.3$ is an efficient technique to select HzRGs. However, in addition to these USS sources, many HzRGs have been discovered with relatively flat spectral index (e.g., ${\alpha}_{\mathrm{1.4GHz}}^{\mathrm{325MHz}}=-0.75$ for J163912.11+405236.5 at $z=4.88$, jarvis_discovery_2009; ${\alpha}_{\mathrm{1.4GHz}}^{\mathrm{325MHz}}=-1.1$ for HSC J083913.17+011308.1 at $z=4.72$, yamashita_wide_2020; see also miley_distant_2008). As a consequence, a sole reliance on USS could lead to a bias in the statistical physical properties of the HzRGs since the diverse nature could be over-represented by USS.
An alternative approach to select HzRGs, which can circumvent the selection bias arising from the uncertain real spectral index distributions, is the Lyman break technique commonly adopted in the search for high-$z$ galaxies. The effectiveness of this technique in discovering HzRGs has been proven by capetti_quest_2025, who identified 39 $g$-dropout HzRGs expected to be at $3<z<4.5$ over an area of $\sim 2,000\mathrm{\ deg^{2}}$, and by yamamoto_wide_2025, who constructed a photometric sample of 146 HzRGs at $z\sim 4$ within a total survey area of $\sim 560\mathrm{\ deg^{2}}$.
The Lyman break color selection technique selects sources that exhibit a clear Lyman break and blue UV continuum in their optical SEDs covered by broadband (ugriz) photometry. Following hildebrandt_cars_2009 that have used a similar filter set, g-dropouts ($3<z<5$) are selected according to
[…]
and r-dropouts ($4<z<6$) are extracted adopting
[…]
@@
We show g- and r-dropout candidates in the upper and lower panels of Figure 17, respectively. The above criteria produce 4,049 valid sources for the r-dropouts, and only 12 of them can be at $z\sim 5$, which is reasonable considering the limited depth of UNIONS. Of $\color[rgb]{0,0,0}63,019$} UNIONS-VLASS sources with reliable BPZ photo-$z$, there are only 2 at $z>4.5$, comparable with the r-dropout selections.
The number of g-dropout-selected sources is $477$, which exceeds the number counts of 53 expected from reliable $3.5<{z}_{\mathrm{photo}}<4.5$ estimates by almost one order of magnitude. This potential excess in the number counts of g-dropouts – if they are not bona fide sources at $z\sim 4$ – could be explained by the contamination of foreground sources such as red elliptical galaxies at intermediate redshifts, whose ${D}_{\mathrm{4000}}$ break due to the absorption of old stellar populations mimic the Lyman break feature. The photometric errors of these foreground contaminations satisfy the selection criteria of the g-dropout (e.g., ono_great_2018). For Subaru HSC, which has comparable optical number counts and similar efficient broadband wavelengths to those of UNIONS, ono_great_2018 found the contamination rate in the $i$-band magnitude bin of 22.5 can reach $\sim 68\%$ and decreases to $\sim 41\%$ at ${i}_{\mathrm{ABmag}}=23.1$ and to $\sim 35\%$ at ${i}_{\mathrm{ABmag}}=23.7$ for g-dropouts. Therefore, even if we only select UNIONS galaxies at the faint end of $i$-band (${i}_{\mathrm{AB}}>23$), the contamination problem can still persist. Similarly, for $r$-dropouts, ono_great_2018 reported contaminate rates of approximately $59\%$ at ${z}_{\mathrm{ABmag}}=22.3$ and around $23\%$ at ${z}_{\mathrm{ABmag}}=23.5$, implying that merely <5<5 $r$-dropouts could be true positives if applying these estimates to UNIONS-VLASS. Nonetheless, under the optimistic assumption of HSC-level contamination rates, our UNIONS-VLASS radio galaxy catalog is expected to yield more than 200 HzRGs at $z\sim 4$, offering strong potential to significantly enhance the statistical robustness of studies on HzRGs.
@@ 5.3 Caveat to the Catalog
The effect of applying $\mathtt{NN\_dist>30^{\prime\prime}}$ to VLASS2 is clearly reflected in the projected linear size, that is, the majority of our RGs are compact (${l}_{\mathrm{s}}\leq 20$ kpc), which means our UNIONS-VLASS RGs have the tendency to be young and evolving populations. Therefore, for HzRGs that are often selected by the USS feature in previous studies, there are only a few of them showing USS at $z>2$ amongst $15,906$ UNIONS-VLASS-FIRST sources with valid ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$. In addition to the $\mathtt{NN\_dist}$ cut, this global trend of being compact for our RGs is more a result of the limited search radius of $1.5^{\prime\prime}$, as shown in Figure 2 and discussed in § 3.1. To maintain the current cleanliness of the UNIONS-VLASS catalog, if $\mathtt{NN\_dist\leq 30^{\prime\prime}}$ is not adopted, the inclusion of bright and extended radio sources in the UNIONS-VLASS leads to an increase of 20,000 in the sample size, which is about 1/3 of the total number of sources with $\mathtt{NN\_dist\leq 30^{\prime\prime}}$ in VLASS2. We, thereby, remind the readers to be cautious about the intrinsic and unresolved bias arising from the design of this work. Nonetheless, the vast size of this catalog still ensures statistical studies of radio AGNs across various evolution stages, offering valuable insights into the connections between radio AGNs and their host galaxies.
@@ 6 Summary
We presented radio color-color (${\alpha}_{\mathrm{1400MHz}}^{\mathrm{144MHz}}-{\alpha}_{\mathrm{3000MHz}}^{\mathrm{1400MHz}}$) based on LoTSS, FIRST, and VLASS observations, and further plotted spectral indices as a function of redshift and the projected linear size of the synchrotron-emitting volume. We found a great diversity of radio classifications, including compact steep spectrum and peaked spectrum sources that are likely young radio galaxies, ultra steep spectrum sources which may have significantly evolved, and inverted and upturned spectrum that may be linked to transient sources. There are $\sim 10,000$ additional radio galaxies corresponding to more than one optical counterpart. These sources are merger candidates that host RLAGNs. The radio classifications/RLAGN evolution and merger-RLAGN connections will be further studied in future works.
@@ Appendix A Radio Morphology of Sources with $\mathtt{NN\_dist}<30^{\prime\prime}$
We randomly choose 36 VLASS radio sources with $\mathtt{NN\_dist}<30^{\prime\prime}$ and show their Epoch 2 QL images in Figure 18.
@@ Appendix C UNIONS-VLASS r-band Source Counts for RL and RQAGNs
Figure 20: Optical $r$-band source counts for 63,019 UNIONS-VLASS radio galaxies (circles). In both panels, squares represent RL and crosses represent RQ populations. Left: the UNIONS-VLASS sources are divided into RL and RQAGNs based on the radio loudness, where RLAGNs have $\log\mathcal{R}_{\mathrm{obs}}>1$. Right: RL and RQAGNs are separated via the 1.4 GHz spectral luminosity, where RLAGNs have ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$.
@@
We further divide RL and RQ populations based on their 1.4 GHz spectral luminosities (see § 4.3 for details). We see that RQAGNs are still the dominant populations for galaxies at ${r}_{\mathrm{ABmag}}<18.5$. However, at ${r}_{\mathrm{ABmag}}>20$, the source counts of RQAGNs do not decrease as steeply as the trend based on $\mathcal{R}_{\mathrm{obs}}$, which is similar to that of radio quasars (see, e.g, yamashita_wide_2018). With the increasing $r$-band magnitude, there is a lower possibility of finding RQAGNs compared to RL populations. In summary, from this aspect, both RL and RQAGNs are hosted by a large variety of galaxy populations.
@@ Appendix D Radio-loud AGNs
Star formation can contribute to the non-thermal emission, and rigorously $\log\mathcal{R}_{\mathrm{obs}}>2$ may be adopted to exclude star-forming/starburst galaxies that mimic RLAGNs (e.g., chiaberge_origin_2011). This reduces our RLAGNs to 109,901 ($\sim 77.0\%$). The spectral luminosity of free-free emission attributed to star formation can be quantified by (condon_radio_1992):
$\left({\frac{{L}_{\mathrm{T}}}{\mathrm{\ W\ Hz^{-1}}}}\right)\sim 5.5\times 10^{20}\left({\frac{\nu}{\mathrm{GHz}}}\right)^{-0.1}\left[\frac{\mathrm{SFR(\mathit{{M}_{\mathrm{star}}}\geq 5\mathrm{M_{\odot}})}}{\mathrm{M_{\odot}}\ \rm yr^{-1}}\right],$ (D1)
@@
$\left({\frac{{L}_{\mathrm{NT}}}{\mathrm{\ W\ Hz^{-1}}}}\right)\sim 5.3\times 10^{21}\left({\frac{\nu}{\mathrm{GHz}}}\right)^{-\alpha}\left[\frac{\mathrm{SFR(\mathit{{M}_{\mathrm{star}}}\geq 5\mathrm{M_{\odot}})}}{\mathrm{M_{\odot}}\ \rm yr^{-1}}\right].$ (D2)
Assuming an upper limit of $\mathrm{SFR}=300\ \mathrm{M_{\odot}}\ \rm yr^{-1}$ for star-forming main sequence galaxies at $z\gtrsim 2$ (popesso_main_2023), the corresponding luminosities are ${L}_{\mathrm{T,SF}}\sim 1.5\times 10^{23}\mathrm{\ W\ Hz^{-1}}$ and ${L}_{\mathrm{NT,SF}}\sim 1.3\times 10^{24}\mathrm{\ W\ Hz^{-1}}$. In Figure 13 for ${L}_{\mathrm{1.4GHz}}$ and ${L}_{\mathrm{3GHz}}$ as a function of $z$, this ${L}_{\mathrm{NT,SF}}$ only selects radio galaxies at $z\lesssim 1$ where the host galaxies should have much smaller values of SFR compared to high-z sources even if they are rarely observed starburst systems. The real contributions of star-forming activities to the observed synchrotron radiation should be insignificant ($\lesssim 1.5\times 10^{22}\mathrm{\ W\ Hz^{-1}}$ assuming $\mathrm{SFR}=40\ \mathrm{M_{\odot}}\ \rm yr^{-1}$ at $z\sim 0.5$). Therefore, $\log\mathcal{R}_{\mathrm{obs}}>1$ can be safely adopted to select RLAGNs for the UNIONS-VLASS sample.
We see correlations between ${L}_{\mathrm{1.4GHz}}-\mathcal{R}_{\mathrm{obs}}$ and $\mathcal{R}_{\mathrm{obs}}-{z}_{\mathrm{photo}}$, and these correlations are likely a propagation of the correlation between ${i}_{\mathrm{ABmag}}-{z}_{\mathrm{photo}}$ shown in the left panel of Figure 21 that reflects the cosmological surface brightness dimming. As a result, high-$z$ sources have higher possibilities of being classified as RLAGNs because of their optical faintness. In the right panel of Figure 21, we plot absolute radio magnitude $M_{t}$ as a function of the observed radio loudness $\mathcal{R}_{\mathrm{obs}}$, where $M_{t}$ is defined as (ivezic_optical_2002)
@@
Figure 21: Left: UNIONS $r$-band photometry as a function of the photometric redshift. Right: Absolute radio magnitude as a function of the observed radio loudness, where the dashed line indicates the RL/RQ demarcation of $M_{t}=-25.9$. The golden dots represent RLAGNs based on $\log\mathcal{R}_{\mathrm{obs}}>1$ and the violet ones are RQAGNs defined by $\log\mathcal{R}_{\mathrm{obs}}<1$.
@@ Appendix E UNIONS2000 versus UNIONS5000 for i-band Photometry
We compare the distributions of the rir-i color between UNIONS2000 and UNIONS5000 for 76,454 UNIONS-VLASS radio galaxies produced in both data releases in Figure 22. The median difference between ${r}_{\mathrm{ABmag}}$ and ${i}_{\mathrm{ABmag}}$ is 0.44 in both UNIONS2000 and UNIONS5000. However, in comparison with UNIONS2000, the distribution of rir-i in UNIONS5000 is more convergent, which is positively attributed to that the UNIONS5000 has been additionally processed using Pan-STARRS DR3 $i$-band data.
@@ Appendix F Catalog columns
[…]
@@
Columns for VLASS (gordon_quick_2021)
@@
Columns for UNIONS (gwyn_unions_2025)
@@
Columns for SDSS DR16 (ahumada_16th_2020)
@@
1==bad; 2==barely acceptable; 3==good (field_quality) (13)
@@
Columns for Pan-STARRS DR1 (chambers_pan-starrs1_2016)
@@
Columns for AllWISE (cutri_vizier_2021)
@@
Columns for unWISE (schlafly_unwise_2019)
@@
Columns for FIRST (helfand_last_2015)
@@
Columns for LoTSS DR2 (shimwell_lofar_2022)
@@
flag_z_photo int -1==no valid photo-$z$;
0==valid $\mathrm{UNIONS5000}_{ugriz}$ photo-$z$; 1==valid $\mathrm{UNIONS5000}_{ugri}$ photo-$z$
Z_B_ref float if flag_z_photo=-1, $\mathrm{UNIONS5000}_{ugriz}$ photo-$z$ used but not reliable;
if flag_z_photo=0, $\mathrm{UNIONS5000}_{ugriz}$ photo-$z$ used;
if flag_z_photo=1, $\mathrm{UNIONS5000}_{ugri}$ photo-$z$ used;